OpenStudy (anonymous):

solve the equation by factoring: n^2-36=0 how would i start this equation ?

jimthompson5910 (jim_thompson5910):

are you sure it's not n^2 - 36

OpenStudy (anonymous):

whoops. yeah thats right

OpenStudy (anonymous):

(n-6) (n+6) binomial rule

jimthompson5910 (jim_thompson5910):

so you would use the difference of squares rule to factor

jimthompson5910 (jim_thompson5910):

a^2 - b^2 = (a-b)(a+b)

OpenStudy (anonymous):

so what you are saying is that the A value is n^2 and the B value is -36 ???

jimthompson5910 (jim_thompson5910):

a^2 = n^2 ----> a = n b^2 = 36 -----> b = 6

OpenStudy (anonymous):

okay i see.